2:46 Cost of goods old Cost of goods sold

Can you send it to me as word?
Answer #1

Answer:

Mariners Corp.
Horizontal analysis
A B C=A-B D=C/B
2017 2016 Increase or decrease Change in %
Sales revenue $        60,000 $      50,000 $                     10,000 20.00%
Cost of Goods Sold $        42,000 $      30,000 $                     12,000 40.00%
Gross Profit $        18,000 $      20,000 $                     (2,000) -10.00%
Selling & Administrative expenses $           9,000 $        5,000 $                       4,000 80.00%
Operating Income $          9,000 $      15,000 $                     (6,000) -40.00%
Interest expenses $           2,000 $        2,000 $                             -   0.00%
Income before taxes $           7,000 $      13,000 $                     (6,000) -46.15%
Income tax $           2,000 $        4,000 $                     (2,000) -50.00%
Net Income $           5,000 $        9,000 $                     (4,000) -44.44%
Vertical Analysis
2016 2017
A B=A/sales C D=C/Sales
Amount % Amount %
Sales revenue $        50,000 100.00% $                     60,000 100.00%
Cost of Goods Sold $        30,000 60.00% $                     42,000 70.00%
Gross Profit $        20,000 40.00% $                     18,000 30.00%
Selling & Administrative expenses $           5,000 10.00% $                       9,000 15.00%
Operating Income $        15,000 30.00% $                       9,000 15.00%
Interest expenses $           2,000 4.00% $                       2,000 3.33%
Income before taxes $        13,000 26.00% $                       7,000 11.67%
Income tax $           4,000 8.00% $                       2,000 3.33%
Net Income $           9,000 18.00% $                       5,000 8.33%

In case of any doubt, please comment.